OpenStudy (anonymous):

How to covert this polar form to rectangular form? r=2sin(3theta).

OpenStudy (asnaseer):

do you know the relationship between rectangular and polar coordinates?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i don't know what to do with the 3theta.

OpenStudy (asnaseer):

ok, so first re-write sin(3theta) in terms of just sin/cos theta

OpenStudy (anonymous):

what do you mean. It is already in just sine and cosine

OpenStudy (asnaseer):

e.g. sin(2a) = 2sin(a)cos(a)

OpenStudy (asnaseer):

you may find this page helpful: http://en.wikipedia.org/wiki/List_of_trigonometric_identities

OpenStudy (asnaseer):

basically you need to reduce the expression down to terms that only involve a single theta. your current expression involves 3theta

OpenStudy (anonymous):

you want me to write them using trig identities?

OpenStudy (asnaseer):

have you been taught trig identities?

OpenStudy (anonymous):

yes

OpenStudy (asnaseer):

so you should be able to re-write sin(3a) interms of terms that only involve powers of sin(a) and cos(a)

OpenStudy (asnaseer):

look for "Double-angle, triple-angle, and half-angle formulae" on that link I gave you

OpenStudy (asnaseer):

it has the identity for sin(3a) listed there

OpenStudy (asnaseer):

if you are interested in how to derive it then you can try expanding sin(a+2a) using the expansion for sin(A+B)

OpenStudy (asnaseer):

does this make sense so far?

OpenStudy (anonymous):

a little

OpenStudy (anonymous):

i'm trying to solve for it

OpenStudy (asnaseer):

using the trig identity for sin(3a) you should end up with:\[r=2\sin(3\theta)=2(3\sin(\theta)-4\sin^3(\theta)))=6\sin(\theta)-8\sin^3(\theta)\]

OpenStudy (anonymous):

how did you get that?

OpenStudy (asnaseer):

did you look at the triple angle formulae on that web page?

OpenStudy (anonymous):

oh.I never knew there was a triple identity.My teacher only taught us double and half

OpenStudy (anonymous):

and others simple ones

OpenStudy (anonymous):

ok. this makes since

OpenStudy (asnaseer):

you can derive it by using:\[\sin(A+B)=\sin(A)\cos(B)+\cos(A)\sin(B)\]and letting:\[A=\theta\]\[B=2\theta\]

OpenStudy (anonymous):

oh that's cool. but i don't have time for that. no offence.

OpenStudy (asnaseer):

and then using the identities for \(\sin(2\theta)\) and \(\cos(2\theta)\) to simplify further

OpenStudy (asnaseer):

np :)

OpenStudy (asnaseer):

ok, the next step is to multiply both sides of the resulting expression by \(\sin(\theta)\) to get:\[r=2\sin(3\theta)=2(3\sin(\theta)-4\sin^3(\theta)))=6\sin(\theta)-8\sin^3(\theta)\]therefore:\[r\sin(\theta)=6\sin^2(\theta)-8\sin^4(\theta)\]

OpenStudy (asnaseer):

do you recognise \(r\sin(\theta)\) here - how does this map to rectangular coordinates?

OpenStudy (anonymous):

rsin(theta)=y

OpenStudy (asnaseer):

perfect!, so now we have:\[y=6\sin^2(\theta)-8\sin^4(\theta)\]

OpenStudy (asnaseer):

next, do you know how \(\sin(\theta)\) is mapped to rectangular coordinates?

OpenStudy (anonymous):

yes. it relates to the unit circle where y is the sin and r is the radius.

OpenStudy (asnaseer):

you already know that:\[r\sin(\theta)=y\]so just divide both sides by r to get:\[\sin(\theta)=\frac{y}{r}\]

OpenStudy (asnaseer):

and hopefully you know the relationship between \(r^2\) and \(x^2\) and \(y^2\)?

OpenStudy (anonymous):

ya. i understand that

OpenStudy (anonymous):

yes\[x ^{2}+y ^{2}=r ^{2}\]

OpenStudy (asnaseer):

perfect, so using this we can get:\[\sin(\theta)=\frac{y}{\sqrt{x^2+y^2}}\]

OpenStudy (asnaseer):

substitute this into the expression and you should end up with your desired answer

OpenStudy (anonymous):

so \[y=6\sin(\theta)-8\sin ^{3}(\theta) \]

OpenStudy (asnaseer):

not quite:\[y=6\sin^2(\theta)-8\sin^4(\theta)\]substitute:\[\sin(\theta)=\frac{y}{\sqrt{x^2+y^2}}\]into this expression

OpenStudy (asnaseer):

to get an expression that just has x and y in it

OpenStudy (anonymous):

i did. r times\[y/\sqrt{ x^{2}+y ^{2}}\] simplifies to y. becasue \sqrt{ x^{2}+y ^{2} can be rewritten as r

OpenStudy (anonymous):

i have a question about before. why did you write rsin instead of just r

OpenStudy (asnaseer):

recall that we simplified our expression down to:\[y=6\sin^2(\theta)-8\sin^4(\theta)\tag{a}\]then we showed that:\[\sin(\theta)=\frac{y}{\sqrt{x^2+y^2}}\tag{b}\]so final step is just to substitute the expression for \(\sin(\theta)\) from equation (b) into equation (a)

OpenStudy (asnaseer):

we could have also got to the answer by not first multiplying by \(\sin(\theta)\) and just working with r. But that would have involved more square roots. This way we have even powers of \(\sin(\theta)\) which simplifies some of the expressions.

OpenStudy (asnaseer):

hint - using equation (b) we get:\[\sin^2(\theta)=\frac{y^2}{x^2+y^2}\]and:\[\sin^4(\theta)=\frac{y^4}{(x^2+y^2)^2}\]

OpenStudy (asnaseer):

so use these to simplify equation (a)

OpenStudy (anonymous):

ok. let me try

OpenStudy (anonymous):

\[rsin=6y/r-8y ^{3}/r ^{3} \] is this right so far?

OpenStudy (asnaseer):

no - we are trying to get to rectangular coordinates

OpenStudy (asnaseer):

you seem to be moving back to polar coordinates

OpenStudy (anonymous):

oh i see my mistake

OpenStudy (asnaseer):